# Role of the ocean for fast atmospheric evolution revealed by machine learning

### v1

@@ Results
Figure 1: Spearman correlation between GraphCast’s T2m error at 24hrs lead time and variables at the sea surface (a) SST for June, (b) SST-T2m for June, (c) SST for December, (d) SST-T2m for December. Hatching indicates where the correlation is not significant according to a two-tailed t-test at the 95% confidence level. Note that the colorbar increases exponentially.
@@
Figure 2: Inferred areas of ocean-driven dynamics, using the technique of lagged correlations between the observed SST anomaly and mean surface heat flux anomaly (a) over June months and (b) over December months. Correlation is shown only for the points that have significant correlation at zero lag, and where the correlation does not change sign between a lead/lag of 72 hours. Red (blue) colouration indicates where there is significant positive (negative) correlation that does not change sign with lagging. Note the colorbar increases exponentially.
@@
Figure 3: Top row: Spearman correlation between GraphCast’s errors and (a) SST climatology for June (b) SST climatology minus T2m climatology for June. Bottom row: Spearman correlation between GraphCast’s errors aggregated to a monthly level and (c) monthly SST and (d) monthly SST-T2m, for June.
@@ Discussion
There are several promising extensions of this approach; it would be interesting to explore how the errors associated with missing sea surface information change as GraphCast’s forecast lead time is increased; since we would expect these errors to grow, there may also be potential to improve GraphCast’s forecast error by a reasonable amount. Given some evidence of GraphCast under-predicting storm intensity (Charlton-Perez et al., 2024), and demonstrations of improvements of storm intensity for a NWP model coupled to an ocean (Mogensen et al., 2017; Vellinga et al., 2020), it would be interesting to explore patterns of SST correlations with storm intensity error. It would also be interesting use this technique to explore interactions between the atmosphere and other slow moving earth system components, such as the land and cryosphere. Given that several other MLWP models, such as Pangu-Weather (Bi et al., 2023) and FourCastNet (Kurth et al., 2023), predict 2-metre temperature without taking sea surface temperature as input, it would be very interesting to explore the similarities and differences in how these model errors relate to sea surface temperature.
@@
We test for correlations being significantly high or low by using a two-sided t-test of the test statistic $r_{a}(n-2)^{1/2}/(1-r_{a}^{2})^{1/2}$ (Von Storch and Zwiers, 2002), where $n$ is the number of data points used to calculate the result. To calculate the significance of the linear model error reduction in Sec. Predicting errors using a simple model, we used bootstrapping in order to calculate a mean and standard error of the AWRMSE for GraphCast and GraphCast corrected by a linear model. The significance is then calculated using a two-sample t test (Von Storch and Zwiers, 2002).
@@
The ERA5 datasets used are available in the Copernicus Climate Change Service, Climate Data Store, (2023), for registered users. The dataset used is the ERA5 hourly data on single levels from 1940 to present (DOI: 10.24381/cds.adbb2d47). Accessed July-August 2025. The GraphCast model code and weights are available via https://github.com/google-deepmind/graphcast. Climatology data is taken from the WeatherBench 2 dataset (Rasp et al., 2024), more details available at https://weatherbench2.readthedocs.io/.

### v2

# arXiv:2602.01904v2[physics.ao-ph] 02 Mar 2026

@@ Results
Figure 1: Spearman correlation between GraphCast’s T2m error at 24hrs lead time and variables at the sea surface. Hatching indicates where the correlation is not significant according to a two-tailed t-test at the 95% confidence level. (a) SST for June, (b) SST-T2m for June, (c) SST for December, (d) SST-T2m for December. Note that the colorbar increases exponentially.
@@
Figure 2: Inferred areas of ocean-driven dynamics, using the technique of lagged correlations between the observed SST anomaly and mean surface heat flux anomaly. Red (blue) colouration indicates where there is significant positive (negative) correlation that does not change sign with lagging, where significance is assessed according to a two-tailed t-test at the 95% confidence level. (a) over June months and (b) over December months. Correlation is shown only for the points that have significant correlation at zero lag, and where the correlation does not change sign between a lead/lag of 72 hours. Note the colorbar increases exponentially.
@@
Figure 3: Top row: Spearman correlation between GraphCast’s errors and sea surface variables. Hatching indicates where the correlation is not significant according to a two-tailed t-test at the 95% confidence level. (a) SST climatology for June (b) SST climatology minus T2m climatology for June. Bottom row: Spearman correlation between GraphCast’s errors aggregated to a monthly level and (c) monthly SST and (d) monthly SST-T2m, for June.
@@ Discussion
There are several promising extensions of this approach; it would be interesting to explore how the errors associated with missing sea surface information change as GraphCast’s forecast lead time is increased; since we would expect these errors to grow, there may also be potential to improve GraphCast’s forecast error by a reasonable amount. Given some evidence of GraphCast under-predicting storm intensity (Charlton-Perez et al., 2024), and demonstrations of improvements of storm intensity for a NWP model coupled to an ocean (Mogensen et al., 2017; Vellinga et al., 2020), it would be interesting to explore patterns of SST correlations with storm intensity error. It would also be interesting use this technique to explore interactions between the atmosphere and other slow moving earth system components, such as the land and cryosphere. Given that several other MLWP models, such as Pangu-Weather (Bi et al., 2023) and FourCastNet (Kurth et al., 2023), predict 2-metre temperature without taking sea surface temperature as input, it would be very interesting to explore the similarities and differences in how these model errors relate to sea surface temperature. Recent work has also demonstrated features in GraphCast’s latent dimensions that relate to sea ice extent and ocean activity (MacMillan and Ouellette, 2025); this technique could complement our approach to reveal further insights about how GraphCast internally represents the sea surface.
@@ Methods
We test for correlations being significantly high or low by using a two-sided t-test of the test statistic $r_{a}(n-2)^{1/2}/(1-r_{a}^{2})^{1/2}$ (Von Storch and Zwiers, 2002), where $n$ is the number of data points used to calculate the result. To calculate the significance of the error reduction from constructing a linear model to predict errors based on sea surface temperature, we use bootstrapping in order to calculate a mean and standard error of the AWRMSE for GraphCast and GraphCast corrected by a linear model. The significance is then calculated using a two-sample t test (Von Storch and Zwiers, 2002).
@@
The processed ERA5 and GraphCast error datasets analysed during the current study are available in the Zenodo repository, https://doi.org/10.5281/zenodo.18470825. The full GraphCast error dataset generated during the current study is not publicly available due to storage limitations, but is available from the corresponding author on reasonable request. The ERA5 datasets used are available in the Copernicus Climate Change Service, Climate Data Store, (2023), for registered users. The dataset used is the ERA5 hourly data on single levels from 1940 to present (DOI: 10.24381/cds.adbb2d47). Accessed July-August 2025. The GraphCast model code and weights are available via https://github.com/google-deepmind/graphcast. Climatology data is taken from the WeatherBench 2 dataset (Rasp et al., 2024), more details available at https://weatherbench2.readthedocs.io/.
Code Availability
The underlying code for this study is available at https://github.com/bobbyantonio/graphcast_ocean_interaction.
@@ Acknowledgements
Author Contributions
All authors contributed to the study conception and design. BA generated the datasets and performed the analysis. The first draft of the manuscript was written by BA and all authors edited subsequent versions of the manuscript.
Competing Interests
All authors declare no financial or non-financial competing interests.